OpenStudy (anonymous):

Simplify the raidcal. 7 radical 72

OpenStudy (anonymous):

42 rad 2

OpenStudy (anonymous):

\[7\sqrt{72}\] \[= 7\sqrt{6*6*2}\] Take out 6² and you'll get \[7*6\sqrt{2}\] \[=42\sqrt{2}\]